# AdaMesh: Personalized Facial Expressions and Head Poses for Adaptive Speech-Driven 3D Facial Animation

### v3

@@ 3.2 Pose Adapter
To overcome the averaged generation problem in pose style adaptation, we propose the pose adapter and formulate the adaptation as a simple but efficient retrieval task instead of fine-tuning modules. As illustrated in Fig. 2, we first learn a discrete latent codebook $\mathcal{Z}=\{\bm{z_{k}}\in\mathbb{R}^{d_{\bm{z}}}\}_{k=1}^{M}$ of head poses with the novel sequence-encoding VQ-VAE (Siyao et al., 2022), and predicts the speech-relevant head poses with Pose GPT. The VQ-VAE training requires no speech input, so various video datasets of a larger amount can be utilized. This design enables the coverage of a wide range of poses, which is well-suited for learning a realistic pose manifold (Ng et al., 2022). A semantic-aware pose style matrix is derived for each pose sample in the training sample to create a database. We retrieve the nearest matrix in the pose style matrix database as the adapted style matrix in the adaptation stage. The simple retrieval strategy for adaptation is based on the assumption that the possible pose styles are covered in the VQ-VAE learning with the large-scale dataset.
@@
Discrete Pose Space. Given a head pose sequence $\bm{\Psi}=[\bm{\psi}_{1},\ldots,\bm{\psi}_{T}]\in\mathbb{R}^{T\times 3}$, the VQ-VAE encoder, which consists of 1D convolutional layers, firstly encodes it into context-aware features $\hat{\bm{Z}}\in\mathbb{R}^{T^{\prime}\times d_{\bm{z}}}$, where $T^{\prime}=T/w$ denotes the frame numbers of the downsampled features. Then, we obtain the quantized sequence $\bm{Z_{q}}$ via an element-wise quantization function $Q$ that maps each element in $\bm{Z_{q}}$ to its closest codebook entry:
$\bm{Z_{q}}=Q(\hat{\bm{Z}}):=(\underset{\bm{z_{k}}\in\mathcal{Z}}{\arg\min}\|\hat{\bm{z}}_{t}-\bm{z}_{k}\|_{2})\in\mathbb{R}^{T^{\prime}\times d_{\bm{z}}}.$ (2)
@@
The quantized sequence is then reconstructed by the VQ-VAE decoder. Three 1D convolutional layers comprise the VQ-VAE decoder.
@@
Pose GPT. With the learned discrete pose space, we build a Pose GPT network Vaswani et al. (2017); Siyao et al. (2022) that maps the speech and pose style embedding into the discrete codes of the target poses. The HuBERT feature and mel spectrogram of input speech are sent into several Transformer layers (Vaswani et al., 2017) as the speech content features. HuBERT feature conveys more linguistic information while mel spectrogram contains prosody information (Chen et al., 2022). We utilize these two features for the reason that head poses are correlated to both the speech semantics and the speech prosody (e.g., stress, tone). We assign each pose sample in the training dataset with a one-hot label as the pose style embedding. It is passed through several linear layers, concatenated with the speech content features, and then fed into the Pose GPT comprising six cross-conditional Transformer layers. The predicted codes are finally sent into the VQ-VAE decoder to obtain the continuous head pose sequence. We train this Pose GPT using the teacher-forcing training scheme with the cross-entropy loss.
@@
Semantic-Aware Pose Style Matrix. Given a speech utterance of $T$ frames and the corresponding head pose sequence, HuBERT (Hsu et al., 2021) assigns each speech frame into one of 512 clusters, while the VQ-VAE encoder transforms the pose sequence into the latent sequence $\hat{\bm{Z}}$. Thus, each frame in $\hat{\bm{Z}}$ is labeled with a cluster which contains the speech semantic information. We calculate the semantic-aware pose style matrix $S\in\mathbb{R}^{512\times d_{z}}$ for each pose sequence as:
$S_{j}=\sum_{i=1}^{T^{\prime}}\hat{\bm{Z}}_{i}\cdot\delta(L_{i},j)/\sum_{i=1}^{T^{\prime}}\delta(L_{i},j),j=[1,2,\dots,512],$ (3)
@@
where $L_{i}$ denotes the cluster label for $i$ frame of $\hat{\bm{Z}}$ and $\delta$ is the Kronecker symbol that outputs 1 when the inputs are equal else 0. This equation aggregates all the poses which have the same semantic unit of HuBERT. Since the clustering step in HuBERT models the semantic information in acoustic units and the VQ-VAE encoder models the temporal pose style information, the derived $S$ represents the pose style that considers semantics.
@@
Procedure of the Pose Adapter. The pretrain, adaptation, and inference procedures of the pose adapter are listed in Alg. 2. Note that the adaptation stage involves no parameter updating.
Algorithm 2 Pretrain, Adaptation, and Inference of Pose Adapter.
@@
1:  Pretrain: Assign each pose sequence in the source training dataset with a one-hot pose style embedding. Train the VQ-VAE and Pose GPT with source training dataset. Calculate the semantic-aware pose style matrix for each sequence in the training dataset and establish the database for all pose style matrices and their corresponding pose style embedding.
@@
2:  Adaptation: Given an unseen pair of reference speech and pose sequence, calculate the semantic-aware pose style matrix for the reference data. Find the nearest pose style matrix in the established database by computing the L1 distance, and adopt the corresponding pose style embedding for inference.
@@
3:  Inference: Given a driving speech, and utilizing the retrieved pose style embedding, the pose adapter generates head poses with the pre-trained Pose GPT and the VQ-VAE decoder.
4 Experiments
@@
Dataset. Due to the scarcity of 3D animation datasets with personalized talking styles, we utilize the 2D audio-visual talking face dataset. For expression adapter pre-train, we utilize Obama weekly footage (Suwajanakorn et al., 2017) with a neutral talking style. For pose adapter pre-train, VoxCeleb2 (Chung et al., 2018) is utilized, which contains about 1M talking-head videos of different celebrities with diverse and natural head poses. In the adaptation stage, we introduce MEAD (Wang et al., 2020) for the expression adapter, since MEAD consists of videos featuring 60 actors with 7 different emotions. We introduce VoxCeleb2 test dataset for the adaptation of the pose adapter.
@@
Dataset Processing. We resample the 2D videos as 25fps and employ the 3D face reconstruction method DECA (Feng et al., 2021) to extract the identity, expression, and head pose. Given these three parameters, we can compute the 3D vertices of the per-frame meshes. For the audio track, we downsample the raw wave as 16000Hz and adopt the base HuBERT model (Hsu et al., 2021) to extract speech features. Since the frame rates of video and audio features are different, we interpolate the audio features and align them to the same length (Cudeiro et al., 2019).
Baseline Methods. We compare AdaMesh with several state-of-the-art methods. 1) FaceFormer (Fan et al., 2022), which first introduces autoregressive Transformer-based architecture for speech-driven 3D facial animation. 2) CodeTalker (Xing et al., 2023), which is the first attempt to model facial motion space with discrete primitives. 3) GeneFace (Ye et al., 2023), which proposes a domain adaptation pipeline to bridge the domain gap between the large corpus and the target person video. It can also generate realistic head poses with a template pose sequence. 4)Imitator (Thambiraja et al., 2023), which learns identity-specific details from a short input video and produces facial expressions matching the identity-specific speaking style. 5)FACIAL (Zhang et al., 2021), which focuses on predicting dynamic face attributes, including head poses. GeneFace and FACIAL are photo-realistic talking face methods, we therefore reconstruct 3D meshes Feng et al. (2021) from the generated videos for comparison.
5 Results
5.1 Quantitative Evaluation
Evaluation Metric. To measure lip synchronization, we calculate the lip vertex error (LVE), which is widely used in previous methods (Fan et al., 2022; Thambiraja et al., 2023). Since MEAD dataset contains significant emotion, we follow Peng et al. (2023) to calculate the emotional vertex error (EVE) to evaluate if the method can capture the personalized emotional talking style. The diversity scores (Siyao et al., 2022; Chen et al., 2023) for expressions (Div-$\bm{\delta}$) and poses (Div-$\bm{\psi}$) are separately computed to evaluate if the methods generate vivid and diverse motions. The landmark standard deviation (LSD) (Zhang et al., 2021; Xing et al., 2023) is used to measure the pose dynamics. The FID score (Siyao et al., 2022; Zhu et al., 2023) for motion sequence is also employed to estimate the pose realism.
Evaluation Results. The results are shown in Tab.1. For facial expressions, it can be observed that our method obtains lower LVE than other methods, which indicates AdaMesh has more accurate lip movements. FaceFormer and CodeTalker have poorer performance, since the LVE score is calculated on MEAD and Obama datasets, and these two methods have poor generalizability to unseen speaker identity. AdaMesh also achieves the lowest EVE. It suggests that our approach can efficiently capture the facial style in the adaptation. A higher diversity score for expressions confirms that our approach generates more vivid facial dynamics. For head poses, FaceFormer, CodeTalker, and Imitator cannot generate head poses. GeneFace incorporates a real head pose sequence for presentation and FACIAL learns over-smoothed poses. The FID scores show that our approach generates more realistic head poses, while LSD and diversity scores confirm our approach generates head poses with higher diversity.
@@ 5.1 Quantitative Evaluation
Table 1: Quantitative evaluation and user study results. User study scores are with 95% confidence intervals. Exp-Richness value reveals the percentages of users who prefer the compared method other than AdaMesh in A/B tests. Value lower than 50% indicates that users prefer AdaMesh.
@@
FaceFormer 4.67 3.46 6.42 1.05 0.0 0.0 4.05$\pm$0.15 1.46$\pm$0.10 0.80%
Codetalker 4.43 3.30 6.48 1.05 0.0 0.0 4.10$\pm$0.14 1.42$\pm$0.11 0.80%
FACIAL 3.44 2.48 6.32 1.68 0.69 1.50 3.88$\pm$0.21 3.85$\pm$0.20 11.7%
GeneFace 3.54 2.47 6.56 1.38 0.70 1.49 4.01$\pm$0.18 3.79$\pm$0.17 12.3%
Imitator 3.15 2.33 5.97 1.05 0.0 0.0 3.82$\pm$0.18 1.42$\pm$0.11 29.8%
Ground Truth 7.85 1.37 3.01 4.30$\pm$0.14 4.12$\pm$0.23 41.3%
AdaMesh (Ours) 2.91 2.25 7.10 0.90 0.87 1.57 4.12$\pm$0.19 4.22$\pm$0.11
5.2 Qualitative Results
Since quantitative metrics cannot fully reflect the perceptual observation, we conduct qualitative evaluation from three aspects, as shown in Fig. 3.
@@ 5.2 Qualitative Results
Speech-Lip Synchronization. To evaluate the lip synchronization performance, we illustrate three keyframes speaking the vowels and consonants. The visual results show that our approach generates more accurate lip movements than other methods. AdaMesh has more pronounced mouth open when speaking /ou/ and /ei/. FaceFormer and CodeTalker close their mouth tighter on /b/ than other methods, even the ground truth, since they are directly trained on the 3D scanned vertices. Other methods are trained or fine-tuned on the reconstructed data from 2D video, which are less precise than the scanned data.
@@
Facial Expression Richness. To evaluate if the methods can capture the facial talking style from the reference video. We compare GeneFace, Imitator, and our AdaMesh on MEAD dataset. It can be observed that our approach preserves the expressive facial style and obviously gains richer facial details, especially around the cheek and lip regions. AdaMesh even gains more expressive details and accurate lip movements than the ground truth. The MoLoRA parameters in AdaMesh act as a plugin of facial style for the pre-trained model, and the pre-trained model on a larger scale data has guaranteed lip synchronization.
@@
Head Pose Naturalness. We adapt our pose adapter on the VoxCeleb2-Test dataset. As shown in Fig. 3(c), we illustrate frames in a sentence that contains evident emotion. A sentence that conveys strong emotions is accompanied by more evident head poses. Our approach produces head movements with a greater range of amplitudes, and the results are closer to the ground truth. Following FACIAL (Zhang et al., 2021), we also plot the landmark tracemaps for a sentence that conveys negative semantics. The head poses with more degrees of freedom confirm the diversity of AdaMesh.
@@
User Study. We conduct a user study to measure the perceptual quality of these methods. Fifteen users with professional English ability participate in this study. For lip synchronization and pose naturalness, we adopt mean-opinion-score (MOS) tests to ask users to score each video from one (worst) to five (best). Our approach has achieved the highest score on these two aspects. For expression richness, we adopt A/B test to ask users to select a better video between AdaMesh and another method. Our approach has significantly outperformed other methods, including the ground truth. This is attributed to the image jittering brought by Feng et al. (2021); Li et al. (2017). AdaMesh has filtered out these disturbances in the pre-training of the expression adapter.
Figure 3: Qualitative comparison with different methods. (a) shows lip movements on Obama dataset with neutral talking style. (b) is for observation of facial expressions on the emotional MEAD dataset. (c) and (d) show head poses and corresponding landmark tracemaps on the VoxCeleb2-Test dataset. The first row displays the words or sentences that these frames are pronouncing.
5.3 Approach Analysis
Ablation Study for Expression Adapter. For expression adapter, we separately remove the style encoder, identity encoder, and MoLoRA to verify the necessity of these modules. 1) As shown in Tab. 3, the removal of the style encoder leads to a drop of vertices errors and diversity metrics by a large margin. It suggests that the style encoder provides essential details which cannot be predicted from speech. 2) The absence of identity also causes a slight performance drop since the expression and identity parameters are not fully disentangled (Li et al., 2017; Feng et al., 2021). The identity parameters also contain person-specific talking style. 3) We fine-tune the full expression adapter instead of updating MoLoRA parameters. It can be observed that the performance significantly decreases since fine-tuning with few data damages the generalization ability of the pre-trained model.
Ablation Study for Pose Adapter. For pose adapter, we test three settings. 1) The retrieval of the pose style matrix is removed by randomly selecting a pose style embedding in the training data for inference. All metrics drop significantly due to the unawareness of the reference’s personalized talking style. 2) w/o HuBERT feature, where only the mel spectrogram is utilized for the speech representation. The LSD and diversity scores show that HuBERT features contain semantic information, which is vital to the diverse head pose generation. 3) After removing VQ-VAE, the model directly predicts the continuous pose values. The LSD and diversity scores drop to near zero. It suggests that VQ-VAE provides realistic quantized codes which is easier to predict. To further investigate the efficacy, we plot the yaw degree of head poses in Fig. 4(c). The pose curve of AdaMesh has the largest variation, which is closer to the ground truth. It confirms the results in quantitative and qualitative experiments. We also utilize t-SNE (Van der Maaten & Hinton, 2008) to visualize the pose style matrix for the selected exciting and peaceful head poses. The visualization result shows that different styles are clustered into different groups. This proves that the semantic-aware pose style matrix contains meaningful pose style information.
Amount of Adaptation Data. We only analyze the impact of the amounts of adaptation data on the performance of the expression adapter, since the pose adapter actually is a zero-shot model with the retrieval strategy. Fig. 4(a) demonstrates the performance vs. adaptation data amount for AdaMesh. Points within the same color family represent the same method, while the same marker corresponds to the computation of the same metric. Thus the points with the same marker can be compared. There are two y-axis in the figure, and the legend next to each y-axis indicates what metric this y-axis represents. We can conclude from this figure that, increasing the amount of adaptation data can obviously improve the performance. AdaMesh achieves significantly better performance than Imitaor with the same amount of adaptation data.
Amount of adaption Parameters. To better estimate the impact of the amounts of MoLoRA parameters on the performance of the expression adapter, we evaluate the performance of MoLoRA with different rank size combinations and the vanilla LoRA (the point in Fig. 4(b) with the maximum number of parameters). The point with the highest diversity score corresponds to the rank sizes of $[4,8,16,32]$. This setting also achieves the lowest LVE and EVE scores. Furthermore, when optimizing roughly the same number of parameters, our approach outperforms Imitator in all metrics.
Table 2: Ablation for expression adapter.

### v4

@@ III-C Pose Adapter
To overcome the averaged generation problem in pose style adaptation, we propose the pose adapter and formulate the adaptation as a simple but efficient retrieval task instead of fine-tuning modules. As illustrated in Fig. 4, we first learn a discrete latent codebook $\mathcal{Z}=\{\boldsymbol{z_{k}}\in\mathbb{R}^{d_{\boldsymbol{z}}}\}_{k=1}^{M}$ for head poses with the novel sequence-encoding VQ-VAE [45], and then predicts the discrete head pose codes from the driving speech with PoseGPT. The VQ-VAE training requires no speech input, so various video datasets of a larger amount can be utilized. This design enables the coverage of a wide range of pose styles, which is well-suited for learning a realistic pose manifold [46]. A semantic-aware pose style matrix is derived for each pose sample in the training sample to create a database. We retrieve the nearest matrix in this database as the adapted style matrix in the adaptation stage. The simple retrieval strategy for adaptation is based on the assumption that the possible pose styles have been covered in the VQ-VAE learning with the large-scale dataset.
@@
Discrete Pose Space. Given head pose sequence in continuous domain $\boldsymbol{\Psi}=[\boldsymbol{\psi}_{1},\ldots,\boldsymbol{\psi}_{T}]\in\mathbb{R}^{T\times 3}$, the VQ-VAE encoder, which consists of 1D convolutional layers, firstly encodes it into context-aware features $\hat{\boldsymbol{Z}}\in\mathbb{R}^{T^{\prime}\times d_{\boldsymbol{z}}}$, where $T^{\prime}=T/w$ denotes the frame numbers of the downsampled features. Then, we obtain the quantized sequence $\boldsymbol{Z_{q}}$ via an element-wise quantization function $Q$ that maps each element in $\hat{\boldsymbol{Z}}$ to its closest codebook entry:
$\boldsymbol{Z_{q}}=Q(\hat{\boldsymbol{Z}}):=(\underset{\boldsymbol{z_{k}}\in\mathcal{Z}}{\arg\min}\|\hat{\boldsymbol{z}}_{t}-\boldsymbol{z}_{k}\|_{2})\in\mathbb{R}^{T^{\prime}\times d_{\boldsymbol{z}}}.$ (2)
@@
$\boldsymbol{Z_{q}}$ is then reconstructed to continuous poses by VQ-VAE decoder, which consists of three 1D convolutional layers.
@@
PoseGPT. With the learned discrete pose space, we build a PoseGPT network [17, 45] that maps the driving speech and pose style embedding into the discrete codes of the target poses. The HuBERT feature and mel spectrogram of the driving speech are utilized for speech representation to consider both speech semantics and speech prosody (e.g., stress, tone) [21] information. We assign each pose sample in the training dataset with a one-hot label as the pose style embedding. The speech representation and pose style embedding are sent into PoseGPT, which consists of six cross-conditional Transformer layers [17]. PoseGPT predicts discrete codes autoregressive by selecting the output with the highest probability at each step. The predicted codes are finally sent into the VQ-VAE decoder to obtain the continuous head pose sequence. We train PoseGPT using the teacher-forcing training scheme with the cross-entropy loss.
@@
Semantic-Aware Pose Style Matrix. Given a speech utterance of $T$ frames and the corresponding head pose sequence, HuBERT [37] assigns each speech frame into one of 512 clusters, while the VQ-VAE encoder transforms the pose sequence into the latent sequence $\hat{\boldsymbol{Z}}$. Thus, each frame in $\hat{\boldsymbol{Z}}$ is labeled with a cluster which contains the speech semantic information. We calculate the semantic-aware pose style matrix $S\in\mathbb{R}^{512\times d_{z}}$ for each pose sequence as:
$S_{j}=\sum_{i=1}^{T^{\prime}}\hat{\boldsymbol{Z}}_{i}\cdot\kappa(L_{i},j)/\sum_{i=1}^{T^{\prime}}\kappa(L_{i},j),j=[1,\dots,512],$ (3)
@@
where $L_{i}$ denotes the cluster label for $i$ frame of $\hat{\boldsymbol{Z}}$ and $\kappa$ is the Kronecker symbol that outputs 1 when the inputs are equal else 0. This equation aggregates all the poses which have the same semantic unit of HuBERT. Since the clustering step in HuBERT [37] captures the semantic information in acoustic units and the VQ-VAE encoder models the temporal pose style information, the derived $S$ represents the pose style that considers semantics.
@@
Procedure of Pose Adapter. The pipeline of the pose adapter is listed in Alg. 2. Note that the adaptation stage involves no parameter updating.
Algorithm 2 Pipeline of Pose Adapter.
@@
  Pretrain: For each pose sequence in the source training dataset, assign it with a one-hot pose style embedding, and calculate its semantic-aware pose style matrix with its corresponding speech. Establish the database for all pose style matrices and their paired pose style embedding. Train the VQ-VAE and PoseGPT with the source training dataset.
@@
  Adaptation: Given an unseen pair of reference speech and pose sequence, calculate the semantic-aware pose style matrix for the reference data. Find the nearest pose style matrix in the established database by computing the L1 distance, and adopt the corresponding pose style embedding for inference.
@@
  Inference: Given a driving speech, and utilizing the retrieved pose style embedding, the pose adapter generates continuous head poses with the pre-trained PoseGPT and the VQ-VAE decoder.
Loss Function. The loss function of VQ-VAE is:
$\mathcal{L}_{VQ}=\mathcal{L}_{rec}(\hat{\boldsymbol{\Psi}},\boldsymbol{\Psi})+\left\|\text{sg}[\hat{\boldsymbol{z}}]-\boldsymbol{z_{q}}\right\|+\gamma\left\|\boldsymbol{z}-\text{sg}[\boldsymbol{z_{q}}]\right\|$ (4)
where $\hat{\boldsymbol{\Psi}}$ and $\boldsymbol{\Psi}$ denote predicted and ground truth pose sequences, and $\text{sg}[\cdot]$ denotes the operation of stop gradient. $\mathcal{L}_{rec}(\hat{\boldsymbol{\Psi}},\boldsymbol{\Psi})$ can be formulated as:
$\mathcal{L}_{\text{rec }}(\hat{\boldsymbol{\Psi}},\boldsymbol{\Psi})$ $=\left\|\hat{\boldsymbol{\Psi}}-\boldsymbol{\Psi}\right\|_{1}+\alpha_{1}\left\|\hat{\boldsymbol{\Psi}}^{\prime}-\boldsymbol{\Psi}^{\prime}\right\|_{1}$ (5)
$+\alpha_{2}\left\|\hat{\boldsymbol{\Psi}}^{\prime\prime}-\boldsymbol{\Psi}^{\prime\prime}\right\|_{1}$
where $\boldsymbol{\Psi}^{\prime}$ and $\boldsymbol{\Psi}^{\prime\prime}$ represent the $1^{st}$-order (velocity) and $2^{nd}$-order (acceleration) partial derivatives of pose sequence $\boldsymbol{\Psi}$ on time. The velocity and acceleration loss items ensure the generation of smooth head poses and prevent abnormal jitters, which is commonly used in previous works [23, 45]. $\alpha_{1}$, $\alpha_{2}$ and $\gamma$ denote trade-off coefficients. The PoseGPT is optimized with the MSE loss. To expedite convergence and align with the inference process, we employ a teacher-forcing strategy to train PoseGPT during the early training and gradually transit to the autoregressive training scheme.
IV Experiments
Figure 5: Qualitative comparison with different methods. (a) shows lip movements on Obama dataset with neutral talking style. (b) is for observation of personalized facial expressions on the emotional MEAD dataset. (c) and (d) show head poses and corresponding landmark tracemaps on the VoxCeleb2-Test dataset. The first row displays the words or sentences that these frames are pronouncing.
@@ IV Experiments
Dataset. Due to the scarcity of 3D animation datasets with personalized talking styles, we utilize the 2D audio-visual talking face datasets. For expression adapter pre-train, we utilize Obama weekly footage [47] with a neutral talking style. For pose adapter pre-train, VoxCeleb2 [48] is utilized, which contains about 1M talking-head videos of 6112 speakers with diverse and natural head poses. In the adaptation stage, we introduce MEAD [49] for the expression adapter, since MEAD consists of videos featuring 60 actors with 7 different emotions. We utilize VoxCeleb2-Test dataset for the adaptation of the pose adapter, and the speakers are unseen in pre-train.
@@
Dataset Processing. We resample 2D videos as 25fps. To derive 3D data from 2D data, we employ the state-of-the-art face reconstruction method, SPECTRE [36], to obtain FLAME coefficients. Scanned data offers more precise details in lip and forehead movement. The average error of extracted 3D data [38, 36], compared to scanned data, is around 1-2mm. Considering the average width and length of a face are approximately 130mm and 200mm, such an error is not easily perceptible. To further enhance the fidelity of the reconstructed 3D data, we optimize SPECTRE network for each video clip, rather than using a pre-trained face model for inference directly. Given the extracted parameters of identity, expression, and head pose, we can compute the 3D vertices of the per-frame meshes. For the audio track, we downsample the raw wave as 16kHz and adopt the base HuBERT model [37] to extract speech features. Since the frame rates of video and audio features are different, we interpolate the audio features and align them to the same length [1].
@@
TABLE I: Quantitative evaluation and user study results. User study scores are with 95% confidence intervals. Exp-Richness value reveals the percentages of users who prefer the compared method other than AdaMesh in A/B tests. Value lower than 50% indicates that users prefer AdaMesh (Ours). Collab: two adapters collaborate.
@@
FaceFormer 4.67 3.46 6.42 1.05 4.05$\pm$0.15 1.46$\pm$0.10 0.80%
Codetalker 4.43 3.30 6.48 1.05 4.10$\pm$0.14 1.42$\pm$0.11 0.80%
FACIAL 3.44 2.48 6.32 1.68 0.69 1.50 30.65 3.88$\pm$0.21 3.85$\pm$0.20 11.7%
GeneFace 3.54 2.47 6.56 1.38 0.70 1.49 34.13 4.01$\pm$0.18 3.79$\pm$0.17 12.3%
SadTalker 4.80 3.62 6.42 1.38 0.53 1.11 62.40 3.88$\pm$0.14 3.82$\pm$0.15 11.5%
Imitator 3.15 2.33 5.97 1.05 3.98$\pm$0.19 1.42$\pm$0.11 29.8%
Mimic 3.03 2.30 6.67 1.05 4.05$\pm$0.18 1.42$\pm$0.11 40.2%
Ground Truth 0 0 7.85 1.37 3.01 4.30$\pm$0.14 4.12$\pm$0.23 41.3%
AdaMesh (Ours) 2.91 2.25 7.10 0.90 0.87 1.57 13.89 4.12$\pm$0.19 4.22$\pm$0.11
AdaMesh (Collab) 2.93 2.25 7.08 1.66 0.55 1.21 57.43 4.10$\pm$0.18 3.55$\pm$0.20 48.3%
Baseline Methods. We compare AdaMesh with several state-of-the-art methods. 1) FaceFormer [3], which first introduces autoregressive Transformer-based architecture for speech-driven 3D facial animation. 2) CodeTalker [4], which is the first attempt to model facial motion space with discrete primitives. 3) FACIAL [14], which focuses on predicting dynamic face attributes, including head poses. 4) GeneFace [8], which proposes a domain adaptation pipeline to bridge the domain gap between the large corpus and the target person video. It can also generate realistic head poses with a template pose sequence. 5) SadTalker [30], which learns the realistic 3D motion coefficient and generalize to various images. 6) Imitator [9], which learns identity-specific details from a short input video and produces facial expressions matching the personalized speaking style. 7) Mimic [6], which learns disentangled representations of the speaking style and content from facial motions by building two latent spaces. FACIAL, GeneFace and SadTalker are photo-realistic talking face methods. We therefore reconstruct 3D meshes [38] from the generated videos for FACIAL and GeneFace, and convert the BFM [50] to FLAME [19] format [^1] https://github.com/TimoBolkart/TF_FLAME for SadTalker.
V Results
V-A Qualitative Results
To reflect the perceptual quality, we conduct qualitative evaluation from three aspects as shown in Fig. 5. Best viewed with the demo video on the AdaMesh project page.
@@ V-A Qualitative Results
Speech-Lip Synchronization. We illustrate three keyframes speaking the vowels and consonants. The visual results show that our approach generates more accurate lip movements than other methods. AdaMesh has more pronounced mouth open when speaking /ou/ and /ei/. FaceFormer and CodeTalker close their mouth tighter on /b/ than other methods, since they are directly trained on the 3D scanned vertices. Other methods are trained or fine-tuned on the reconstructed data from 2D video, which are less precise than the scanned data.
@@
Facial Expression Richness. From Fig. 5(b), it can be observed that our approach preserves the expressive facial style and obviously gains richer facial details, especially around the cheek and lip regions. AdaMesh even gains more expressive details and accurate lip movements than the ground truth in the supplementary video. The MoLoRA parameters in AdaMesh act as a plugin of facial style for the pre-trained model, and the pre-trained model on the Obama dataset has guaranteed lip synchronization. It can be observed that the introduction of MoLoRA significantly enhances the richness of facial expressions without compromising the accuracy of lip movements. Since the audio-lip sync performance of the adaptation data is typically close to the average of the pre-trained dataset, the MoLoRA strategy focuses on capturing the differences, particularly the expression style, between the adaptation dataset and the pre-trained dataset.
@@
Head Pose Naturalness. As shown in Fig. 5(c), we illustrate frames in a sentence that contains evident emotion. A sentence that conveys negative semantics is accompanied by more evident head poses, which is consistent with the FSD result in the quantitative evaluation. Our approach produces head movements with a greater range of amplitudes, and the results are closer to the ground truth. Following previous works [14, 30], we also plot the landmark tracemaps for a sentence that conveys negative semantics. The head poses with more degrees of freedom confirm the diversity generation of AdaMesh.
@@
User Study. We conduct a user study to measure the perceptual quality of these methods. Fifteen users with professional English ability participate in this study. For lip synchronization and pose naturalness, we adopt mean-opinion-score (MOS) tests to ask users to score each video from one (worst) to five (best). Our approach has achieved the highest scores on these two aspects. For expression richness, we adopt A/B tests to ask users to select a better video between AdaMesh and another method. Our approach has significantly outperformed other methods, including the ground truth. This is attributed to the image jittering brought by [38, 19]. AdaMesh has filtered out these disturbances in the pre-training of the expression adapter.
TABLE II: Comparsion results on VOCASET. FT denotes finetuning with few VOCASET data.
Method LVE (mm) $\downarrow$ Trained on VOCASET
FaceFormer 4.3011 Yes
CodeTalker 4.2902 Yes
Ours w/ FT 4.2237 30s adaptation
Ours wo/ FT 4.2933 No
V-B Quantitative Evaluation
Evaluation Metric. To measure lip synchronization, we calculate the lip vertex error (LVE), which is widely used in previous methods [3, 9]. We follow [22] to calculate the emotional vertex error (EVE) to evaluate if the method can capture the personalized emotional talking style on MEAD dataset [49]. The diversity scores [45, 24] for expressions (Div-$\boldsymbol{\delta}$) and poses (Div-$\boldsymbol{\psi}$) are separately computed to evaluate if the methods generate vivid and diverse motions. The landmark standard deviation (LSD) [14, 4] measures the pose dynamics. The FID score [45, 25] estimates the head pose realism. To evaluate if the head poses are semantically aligned with the speech, we calculate the FSD score by computing the FID scores which correspond to the same HuBERT units and then averaging these scores [37].
Evaluation Results. The results are shown in Tab.I. For facial expressions, it can be observed that our approach obtains lower LVE than other methods, which indicates AdaMesh has more accurate lip movements. FaceFormer and CodeTalker have poorer performance, since the LVE score is calculated on MEAD and Obama datasets, and these two methods have poor generalization ability to unseen speaker identity. AdaMesh also achieves the lowest EVE. It suggests that our approach can efficiently capture the facial style in the adaptation. A higher diversity score for expressions confirms that our approach generates more vivid facial dynamics. For head poses, FaceFormer, CodeTalker, and Imitator cannot generate head poses. GeneFace incorporates a real head pose sequence for presentation and  FACIAL learns over-smoothed poses. The FID scores show that our approach generates more realistic head poses, while LSD and diversity scores confirm our approach generates head poses with higher diversity. AdaMesh also gains the lowest FSD score, suggesting that the generated head poses are semantic aware. To further explore whether the collaboration of two adapters can collectively contribute to the task, we combine the expression and pose adapters, referring it as AdaMesh (Collab). It adopts the MoLoRA strategy to capture speaking styles, and simultaneously generates expression and pose parameters. Compared with AdaMesh (Ours), AdaMesh (Collab) shows a slight decrease in performance related to expression and a significant decrease in pose-related metrics. This is ascribed to the collaborative generation disregarding the intrinsic data characteristics of head poses. It substantiates the rationale for employing separate adapters to capture speaking styles.
TABLE III: Ablation study for expression adapter.